How to calculate how much room temperature water (23.3°C) will be needed to drop the temperature of warm water from 50°C to 42°C.


1
Expert's answer
2021-01-12T07:00:37-0500

Δ1 = 50 – 42 = 80C – cooling of warm water

Q1 = cm1Δ1 = 8cm1

Δ2 = 42 – 23.3 = 18.70C – heating up of room temperature water

Q2 = cm2Δ2 = 18.7cm2

Q1 = Q2

8cm1 = 18.7cm2

8m1 = 18.7m2

m1 = 18.7m2 / 8 = 2.34m2

So, the mass of colder water mast be 2.34 times greater that the mass of warmer water.